OpenStudy (anonymous):

Which operation would you use on the exponents to simplify the expression (9^4)^2? Multiplication Division Addition Subtraction

OpenStudy (anonymous):

(a^b)^c = a^(bc)

OpenStudy (jdoe0001):

\(\large\bf (9^4)^2\implies 9^{4\cdot 2}\)